OpenStudy (anonymous):

A seller has claimed that less than 60 percent of her products are destroyed before 10 years. What is the null hypothesis H0 for testing this claim?

OpenStudy (anonymous):

\[a) H _{0} : \pi < 0.60\] \[a) H _{0} : \pi \ge 0.60\] \[a) H _{0} : \pi \le 0.60\] \[a) H _{0} : \pi > 0.60\]